OpenStudy (anonymous):

Use the Laplace Transform to solve the "integral equation" y'(t)=sin(t)+the integral (from 0 to t) y(t-u)cos(u)du with y(0)=0

OpenStudy (anonymous):

Use the convolution theorem for the transform of the integral. \[y'(t)=\sin t+\int_0^t (t-u)\cos u~du\] where you would have \[\mathcal{L}\left\{\int_0^t f(t-u)g(u)~du\right\}=\mathcal{L}\{f(t)\}\times\mathcal{L}\{g(t)\}\]

OpenStudy (anonymous):

In this case, \(f(t-u)=t-u\) and \(g(u)=\cos u\), so you would have \[\mathcal{L}\left\{\int_0^t (t-u)\cos u~du\right\}=\mathcal{L}\{t\}\times\mathcal{L}\{\cos t\}\]

OpenStudy (anonymous):

there is a y under the integral too. will that change the process?

OpenStudy (anonymous):

Sorry, must have missed it. The question is \[y'(t)=\sin t+\int_0^t y(t-u)\cos u~du\] right? In this case, you can still use the convolution theorem, but set \(f(t-u)=y(t-u)\) (assuming that's written to mean \(y\) as a function of \(t-u\), and not the function \(y(t)\) times \(t-u\)), so you have \[\begin{align*} \mathcal{L}\left\{y'(t)\right\}&=\mathcal{L}\left\{\sin t\right\}+\mathcal{L}\left\{\int_0^t y(t-u)\cos u~du\right\}\\\\ s\mathcal{L}\left\{y(t)\right\}-y(0)&=\mathcal{L}\left\{\sin t\right\}+\mathcal{L}\left\{y(t)\right\}\mathcal{L}\left\{\cos t\right\}\\\\ \mathcal{L}\left\{y(t)\right\}\left(s-\mathcal{L}\left\{\cos t\right\}\right)&=\mathcal{L}\left\{\sin t\right\} \end{align*}\]

OpenStudy (anonymous):

ok i follow you. but this doesn't solve it does it? you would have the laplace of sin(t) divided by (s-laplace cos(t)).....that gives the laplace of y?

OpenStudy (anonymous):

Right, you have to take the transforms of the sine and cosine, and you have an equaiton of the form \(\mathcal{L}\{y\}=\cdots\), then you'd solve for \(y\) by taking the inverse transform, \(\mathcal{L}^{-1}\left\{\mathcal{L}\{y\}\right\}=y=\mathcal{L}^{-1}\{\cdots\}\).

OpenStudy (anonymous):

so I have L(y)= 1/s^3......so to take the inverse laplace transform of each side my final answer would be y=...something. wolfram showed me that the inverse laplace transform of 1/s^3 is t^2/2, but I have no idea how to actually find that. Can you show me how they got that answer?

OpenStudy (anonymous):

wait, i am watching a khan academy video and I think since t^n=n!/s^n+1

OpenStudy (anonymous):

It comes from the fact that \[\mathcal{L}\{t^n\}=\frac{n!}{s^{n+1}}\]

OpenStudy (anonymous):

then 1/s^3 is the same as 2!/2s^2+1

OpenStudy (anonymous):

which gives t^2/2 as the original function?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

YESSSSSSS. so wait, is this all I have to do? I'm done?

OpenStudy (anonymous):

Yes that's it :) The nice thing about solving ODES with initial conditions and the Laplace transform is no pesky \(C\)s to deal with.

OpenStudy (anonymous):

So seriously I am done with this problem? That seemed so simple now that you showed the steps. What the heck lol! Thanks so much

OpenStudy (anonymous):

Yes, you're done! You're welcome!

OpenStudy (anonymous):

Would you be able to help me with something else really quick? It's another laplace transform. All i have to do is find the laplace transform of (1-e^-t)/t but this one isn't as nice as the last one in my opinion.

OpenStudy (anonymous):

Wolfram says the laplace transform of that is ln(1+(1/s)) but I cannot figure out how they got it

OpenStudy (anonymous):

Sure. For this one, you'll use the formula \[\mathcal{L}\left\{\frac{1}{t}f(t)\right\}=\int_s^\infty F(u)~du\] where \(F(u)=\mathcal{L}\{f(t)\}\). In this case, \(f(t)=1-e^{-t}\), so the transform would be \[\mathcal{L}\left\{\frac{1}{t}(1-e^{-t})\right\}=\int_s^\infty \mathcal{L}_s\{1-e^{-t}\}~du\] (The subscript on the Laplace operator is meant to denote that we're transforming from the \(t\) domain to \(s\).

OpenStudy (anonymous):

Also, the formula was taken from this neat resource here: http://tutorial.math.lamar.edu/pdf/Laplace_Table.pdf

OpenStudy (anonymous):

Slight correction, that subscript should be a \(u\), not an \(s\).

OpenStudy (anonymous):

sweet! I think my problem is the 1-e^-t. if it was just the e part the it would just be 1/s+1 but thats not correct is it

OpenStudy (anonymous):

\[\mathcal{L}\left\{1-e^{-t}\right\}=\mathcal{L}\{1\}-\mathcal{L}\{e^{-t}\}=\frac{1}{s}-\frac{1}{s+1}\]

OpenStudy (anonymous):

which when integrated becomes..... ln(u)-ln(u+1) from s to infinity?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

Hmm that integral doesn't seem to converge...

OpenStudy (anonymous):

so...do I take the limit of that then?

OpenStudy (anonymous):

yeah something's not right

OpenStudy (anonymous):

I have a class in 15 min, but I'll get back to this when I have a chance. (unless someone else can clear up the issue here)

OpenStudy (anonymous):

thank you so much! I will wait for your help, you have helped so much in the past!!!

OpenStudy (anonymous):

Alright, it doesn't look like we did anything wrong. I must have been thinking too far ahead of my actual computation. You have \[y=\int_s^\infty\left(\frac{1}{u}-\frac{1}{u+1}\right)~du\] Like you mentioned, we'll need to consider the limit: \[y=\lim_{t\to\infty}\int_s^t\left(\frac{1}{u}-\frac{1}{u+1}\right)~du\] which gives \[y=\lim_{t\to\infty}\bigg[\ln u-\ln (u+1)\bigg]_s^t\] By the FTC we get \[y=\lim_{t\to\infty}\bigg[\ln t-\ln (t+1)\bigg]-\ln s+\ln(s+1)=\ln\frac{s+1}{s}\] as desired.

OpenStudy (anonymous):

this looks great! thanks so much!!!! but what is the FTC?

OpenStudy (anonymous):

Fundamental theorem of calc, the one that says if \(F(x)\) is the antiderivative of \(f(x)\), then \[\int_a^b f(x)~dx=F(b)-F(a)\]